Graph the line with a slope of 2/3 that contains the point (3, –7).
Answer by mananth(16949)

m= 2/ 3

Plug value of the slope and point (3 ,-7 ) in
Y = m x + b
-7.00 = 2 + b
b= -7 - 2
b= -9
So the equation will be
Y = 2/3 x -9
when x=0, y=-9 (0,-9)
when x=3,y=-7 (3,-7)
plot the points and draw the graph

m.ananth@hotmail.ca


m.ananth@hotmail.ca